OpenStudy (anonymous):

Brayden is investing $1,617 in three different stocks: Zycodec, Unix, and Xenon. He will invest 40% of the total between Zycodec and Xenon in the Unix stock. Twice as much as the investment in Zycodec will be given to Xenon. How much will Brayden invest in the Unix stock?

OpenStudy (anonymous):

z stands for zycodec x stands for xenon u stands for unix z + x + u = 1617 .40(x + z) = u 2z = x .40(2z + z) = u .40(3z) = u 1.2z = u now sub 1.2z in for u and 2z in for x z + x + u = 1617 z + (2z) + (1.2z) = 1617 4.2z = 1617 z = 1617/4.2 z = 385 2z = x 2(385) = x 770 = x 1.2z = u 1.2(385) = u 462 = u check... z + x + u = 1617 385 + 770 + 462 = 1617 1617 = 1617 (correct) or he will invest 40% of total between z and x in u .40 (z + x) = u .40(385 + 770) = 462 .40(1155) = 462 462 = 462 (correct) or twice as much as z = x 2z = x 2(385) = 770 770 = 770 (correct) Brayden will invest $ 462 in the Unix stock.

OpenStudy (anonymous):

Thank you, could you help me with this one also?

OpenStudy (anonymous):

I cannot open that file on my computer :/ I don't suppose you could write out the question ?

OpenStudy (anonymous):